Maximum range of a gun is 3920m. What is the velocity of bullet?

Grade:11

2 Answers

Arun
4 years ago
Dear Wasim
 
Maximum range = u^2 /g
 
Now u^2 / g = 3920
 
u^2 = 38416
 
u = 196 m/sec
 
Regards
Arun (askIITians forum expert)
Khimraj
4 years ago
Range = u2sin2\Theta/g
it will be maximum when sin2\Theta =1 
max range = u2/g = 3920
u = 196m/s
Hope it clears.